If you react .994 grams of iron (III) oxide with aluminum, what mass of iron (Fe) will be produced?

                                Fe2O3   +   Al   ®   Al2O3  +  F


1
Expert's answer
2021-03-10T05:59:27-0500

Fe2O3 + 2Al = Al2O3 + 2Fe

The mass of iron produced equals:

m(Fe) = m(Fe2O3) × 2 × Mr(Fe) / Mr(Fe2O3) = 0.994 g × 2 × 55.845 g/mol / 159.69 g/mol = 0.695 g


Answer: 0.695 g